# A FLASH on Blazars

### v2

@@ 3.2 Radio SED-fitting
We performed the SED-fitting using RADIOSED (Kerrison et al. 2024), a software designed to model and fit radio SEDs of astrophysical sources (Figure 4, Appendix LABEL:sec:appendix). This tool is particularly suited for analysing multi-frequency radio data, allowing us to disentangle the contributions of various emission mechanisms by testing multiple spectral models.
@@
Figure 4: Spectral Energy Distribution (SED) fitting for one of the selected blazars in our sample. The models incorporate GLEAM-X, RACS, AT20G, and ALMACAL data. The best-fit model is shown as a solid line, while individual photometric points are marked.
3.2.1 SED-fitting and model selection
@@
RADIOSED supports different functional forms to describe radio SEDs. Throughout this paper, we classify the spectral shapes as follows, consistently with Kerrison et al. (2024):
  • Power Law: A simple model given by
@@
  • Peaked Spectrum: Sources exhibiting a turnover in their SED, where the flux density increases up to a peak frequency $\nu_{\text{peak}}$ and then decreases. This behaviour can indicate synchrotron self-absorption or a young radio source. We model these sources using the functional form (Snellen et al. 1998; Callingham et al. 2017; Slob et al. 2022):
@@
  • Curved Peaked Spectrum: A modification of the peaked spectrum including a curvature term (Orienti et al. 2007, 2010):
@@
  • Re-triggered Peaked Spectrum: A peaked SED with a secondary low-frequency upturn, suggesting episodic jet activity or restarted accretion (Edwards & Tingay 2004). It is modeled as:
    $S_{\nu}=\frac{S_{p}}{1-e^{-1}}[1-e^{(\nu/\nu_{p})^{\alpha_{\mathrm{thin}}-\alpha_{\mathrm{thick}}}}]\Big{(}\frac{\nu}{\nu_{p}}\Big{)}^{\alpha_{\mathrm{thick}}}+S_{0}\nu^{\alpha}.$ (4)
Each model is fitted independently, and the best-fit model is determined through Bayesian model selection. RadioSED computes the Bayes factor to compare the likelihoods of different models, selecting the one with the highest posterior probability.
We adopt non-informative priors (log-uniform/linear-uniform) as described in Table 2 of Kerrison et al. (2024), to ensure a statistically robust classification of spectral types, providing a reliable characterization of the blazar SEDs in our sample.
However, the SED modeling does not account for intrinsic variability between epochs, which can affect the interpretation of blazar spectra. This represents a limitation of our modeling approach that should be considered when analyzing the results.
@@ 4 Results
4.1 Results from the SED-fitting
From the SED fitting of the radio emission, we retrieved the spectral behaviour of the sources, finding that most of them are well described by the Re-triggered Peaked Spectrum, as defined by Equation 4. This smooth behaviour, in contrast to a flat spectrum, suggests that the emission is dominated by a single emitting region rather than by the superposition of multiple compact, self-absorbed regions, as indicated by the classic blazar paradigm (Massardi et al. 2016).
The main parameters retrieved from the SED-fitting (Table 2, Figure 4) are:
  • $\nu_{\mathrm{trough}}$: the frequency corresponding to the minimum of the SED. After this point, if present, the SED starts rising.
  • $\alpha_{\mathrm{trough}}$: the spectral index at frequencies $\nu<\nu_{\mathrm{trough}}$, usually corresponding to GLEAM-X photometric points for our sample.
  • $\nu_{\mathrm{peak}}$: the frequency where, if present, the peak of the SED takes place.
  • $\alpha_{\mathrm{thick}}$: the spectral index at frequencies $\nu_{\mathrm{trough}}<\nu<\nu_{\mathrm{peak}}$, corresponding to the thick part of the peaked component.
  • $\alpha_{\mathrm{thin}}$: the spectral index at frequencies $\nu>\nu_{\mathrm{peak}}$, corresponding to the thin part of the peaked component.
Using the advanced surveys employed in this study, we conclude that at GLEAM-X wavelengths the spectral behaviour is not completely flat ($\alpha\sim-0.5$, Table 2). This finding opens the path to an additional extended, steep-spectrum emission, previously missed (Massaro et al. 2013a, b; Nori et al. 2014). GLEAM-X is, in fact, sensitive to diffuse, low-frequency emission likely from older radio lobes that are missed in higher-frequency surveys.
Figure 5: Relation between the trough frequency $\nu_{\text{trough}}$ and the spectral index at lower frequencies $\alpha_{\text{trough}}$. The colour scale represents the peak frequency $\nu_{\text{peak}}$, which can be used as a proxy for the cooling time of the source. The colours of the markers’ edges represent the three blazar populations: FSRQ, BL-Lac and BCU.
Figure 6: Radio colour-colour diagram of the blazars in our sample, based on the spectral indices retrieved from the SED fitting. The colours represent the three blazar populations: FSRQ, BL-Lac and BCU.
@@ 4.1 Results from the SED-fitting
For most of the sources, the SED extends smoothly towards higher frequencies with a peaked spectrum emission, where it is possible to distinguish a thick part ($\alpha_{\mathrm{thick}}$) and a thin part ($\alpha_{\mathrm{thin}}$).
The median $\nu_{\mathrm{peakrest}}$ is $\sim$ 4.6 GHz and no hint of breaks due to ageing is present at high frequency.
In the radio colour-colour diagram, displayed in Figure 6, it emerges that most of the objects ($\sim$ 57.6 %) are in the fourth quadrant, i.e. have a hint of a peaked spectrum and $\sim$ 10.2% have a flat spectrum, after $\nu_{\mathrm{trough}}$. This is a rather small percentage, indicating that in the GHz domain as well we would expect a single emitting region, rather than the superposition of a number of compact regions of different sizes, self-absorbed at different frequencies.
This single component dominating at $\nu>\nu_{\mathrm{trough}}$ could represent the compact self-absorbed segment that was re-activated at a later time. Meanwhile, the low-energy segment of the SED could delineate an extended remnant from earlier activity (Figure 7). Our results suggest that resolution differences have a limited impact on the main trends observed in the radio SEDs, as most sources are unresolved at GHz frequencies. However, deeper, high-resolution low-frequency surveys (e.g., LOFAR, SKA-Low) will be valuable for disentangling diffuse extended emission from compact jet components.
On the radio colour-colour diagram, BL-Lacs display a wide range of behaviours, whereas FSRQs tend to cluster in the flat region. This indicates that FSRQs undergo highly efficient electron pumping, which significantly boosts high-energy radio synchrotron emissions.
Conversely, both populations seem indistinguishable when considering $\nu_{\mathrm{trough}}$ and $\alpha_{\mathrm{trough}}$ (Figure 6) This may suggest that the extremely low radio frequency emissions are remnants of the source’s history, remaining unchanged by subsequent activities detected at higher frequencies, as the former emissions are confined to the outskirts, while the latter is linked to the nuclear core region.
@@ 5.1 The Radio-gamma correlation
Figure 8: Radio luminosity at 150 MHz (left), 855 MHz (centre) and 20 GHz (right) as a function of the $\gamma$-ray luminosity from Fermi-LAT. The black dashed line and the green line indicate the bisector and the median trend, respectively. The colours indicate the blazar class from the 4FGL catalogue.
Table 3: Median and correlation between the ratio of radio and X-ray luminosities for different object classes (All, FSRQ, BL-Lac). The values are expressed as a function of the luminosities at 150 MHz, 855 MHz, and 20 GHz, relative to the $\gamma$-ray luminosity.
All FSRQ BL-Lac
@@
[…]
Using the redshift estimates, we have drawn a picture of the energy budget spent by the sources in the radio bands from $\sim$ 72 MHz to $\sim$ 30 GHz and in the Fermi-LAT bands. This, combined with the radio SED-fitting described in Section 4.1, allowed us to find the primary emitting components of the object and its evolutionary stage. The median value of the radio luminosity at 855 MHz is 2.5×$\times 10^{26}\mathrm{WHz^{-1}}$ for the spectroscopic sample, with 25% of them having a luminosity $>1.8\times 10^{27}\mathrm{WHz^{-1}}$. 25% of the sources have a $\gamma$-ray luminosity higher than $>1.8\times 10^{27}\mathrm{WHz^{-1}}$ and the median value is $4.3\times 10^{26}\mathrm{WHz^{-1}}$.
The correlation between radio and $\gamma$-ray emission, if present, supports the idea that they are both dominated by the relativistic jet (Giroletti et al. 2016; Mahony et al. 2016). In this work, we can observe how this relation behaves in different radio bands (Figure 8, Table 3). This is particularly useful, since the dominant emitting component varies along the radio spectrum, as we found through the SED-fitting. The luminosity at 855 MHz is closely tied to the $\gamma$-ray luminosity, indicating that, at this frequency, the jet significantly impacts both radio and gamma energy emissions, with a tight relation (correlation coefficient $\rho\sim 0.70$, see Figure 8). This suggests that at 885 MHz, we primarily observe the most powerful segment of the jet, nearest to the nucleus (Figure 7), potentially representing the re-activated section. This is supported by the way the radio-$gamma$-ray relation appears at 20GHz. Consequently, at high radio frequencies, the radio emission is overwhelmingly driven by a strong jet.
In Figure 8 the correlation between 150 MHz and $\gamma$ emission is displayed. Although in this case as well there is a correlation between high and low energy emissions (correlation coefficient $\sim 0.70$), there is an evident offset. A potential explanation might be that the extended synchrotron emission, which is not efficient enough at generating $\gamma$-rays, dominates this part of the spectrum. This scenario is in full agreement with the results obtained from the SED-fitting (Section 4.1).
The FSRQs dominate the bright end of the luminosity in both radio and $\gamma$ and the e Radio-Gamma Correlation keeps holding for both FSRQs and BL-Lacs. Nevertheless, as shown in Table 3, the radio-gamma correlation assumes higher values for FSRQs both at 150 and 855 MHz, approaching the point in which the the radio and $\gamma$ emission are perfectly balanced at lower frequencies. Thus, the FSRQs could saturate the relation already at lower frequencies. This suggests that FSRQs might generally exhibit more effective high-energy pumping across various scales and reflects the presence of a flat spectrum at $\nu>\nu_{trough}$, i.e. the emitting mechanism at 885 and 20 GHz is almost the same.
5.2 The Radio-X-ray correlation
Figure 9: Radio luminosity at 150 MHz (left), 855 MHz (centre) and 20 GHz (right) as a function of the X-ray luminosity from eROSITA. The black dashed line indicates the bisector. The colours indicate the blazar class from the 4FGL catalogue.
To further explore the nature of blazars, we cross-matched our sample with the eROSITA all-sky survey (Merloni et al. 2024). 48 of the sources in the spectroscopic sample have an eROSITA counterpart within 10”. The eROSITA telescope array aboard the Spektrum Roentgen Gamma (SRG) satellite began surveying the sky at 0.2-10 keV. The eROSITA DR1 covers $\sim$ 14056 deg2 with a median sensitivity in the 0.2-2 keV of $\sim$ 10-14erg cm-2s-1. For the cross-matched sources, we were able to study the relationship between X-ray and radio emission. We considered the energy range 0.2 - 2.0 keV for the X-ray luminosity.
Within this spectrum, the sources being analysed exhibit a median luminosity of 6.9×$\times 10^{20}\mathrm{WHz^{-1}}$, with 25% of these sources having a luminosity exceeding $2.8\times 10^{21}\mathrm{WHz^{-1}}$.
Table 4: Median and correlation between the ratio of radio and X-ray luminosities for different object classes (All, FSRQ, BL-Lac). The values are expressed as a function of the luminosities at 150 MHz, 855 MHz, and 20 GHz, relative to the X-ray luminosity in the 0.2-2.0 keV range.
All FSRQ BL-Lac
@@ 5.2 The Radio-X-ray correlation
[…]
The main driver of the emission in this range is the hot corona, which closely traces the accretion disk, especially in FSRQs. The Radio-X-ray correlation is shown in Figure 9. We show the results in Table 4 and we found almost no correlation for the 855 MHz, but a tight one at 150 MHz. In this particular situation, even giving just a glance at Figure 9, it is clear that the global behaviour might be misleading and not truly representative of the physical processes taking place. It is only distinguishing the different classes of blazars that the real picture pops out, as the degree of correlation is clearly unbalanced between FSRQs and BL-Lacs, indicating that the radio and soft X-ray emissions are tightly correlated for FSRQs, whereas in BL-Lacs, the relationship is highly scattered.
The results obtained (4) confirm that FSRQs are in an accretion-driven regime, which fuels both the radio jet and the disk emission. On the other hand, in BL-Lacs the jets are mainly rotation-driven. However, alternative factors could contribute to the observed trends. For instance, BL-Lacs exhibit higher variability in the X-ray band, which may introduce additional scatter in the correlation (Ghisellini et al. 2017).
Moreover, the impact of the circumnuclear medium could play a role in shaping the X-ray emission properties, particularly in FSRQs, where disk-related processes contribute significantly to the total luminosity. A deeper understanding of these effects will require time-domain studies incorporating simultaneous X-ray and radio observations.
When comparing the X-ray luminosity with the 20 GHz emission, it is essential to account for selection bias. The sources in our sample that have both a 20 GHz and a X-ray counterpart represent a highly efficient subset of our data. Consequently, the tighter correlation between X-ray and radio emissions is not surprising, given that the fraction of FSRQs is higher at these frequencies compared to lower frequencies. Deeper millimeter observations are required to gain a comprehensive scientific understanding of the behaviour at higher frequencies.

### v3

@@ 3.2 Radio SED fitting
The SED fitting procedure included all the sources with at least five photometric detections in the radio bands. We performed the SED fitting using RADIOSED (Kerrison et al. 2024), software designed to model and fit radio SEDs of astrophysical sources (Figure 4, Appendix LABEL:sec:appendix). This tool is particularly suited to analysing multi-frequency radio data, thus allowing us to disentangle the contributions of various emission mechanisms by testing multiple spectral models.
@@
Figure 4: Spectral energy distribution fitting for one of the selected blazars in our sample. The models incorporate GLEAM-X, RACS, AT20G, and ALMACAL data. The dark red line is the best fit from the posterior of the selected model type, while individual photometric points are marked. The light red lines are random draws from the posterior representing the range of best-fitting parameters.
@@
RADIOSED supports different functional forms to describe radio SEDs. For this paper and consistent with Kerrison et al. (2024), we classified the spectral shapes as follows:
  • Power law: This class corresponds to a simple model of the form
@@
  • Peaked spectrum: This class includes sources that exhibit a turnover in their SED, where the flux density increases up to a peak frequency, $\nu_{p}$, and then decreases. This behaviour can indicate synchrotron self-absorption or a young radio source. We modelled these sources using the functional form (Snellen et al. 1998; Callingham et al. 2017; Slob et al. 2022)
@@
  • Curved peaked spectrum: This class represents a modification of the peaked spectrum that includes an additional curvature term to better capture broad or asymmetric spectral shapes (Orienti et al. 2007, 2010):
@@
  • Re-triggered peaked spectrum: This class corresponds to a peaked SED with a secondary low-frequency upturn, suggesting episodic jet activity or restarted accretion (Edwards & Tingay 2004). It is modelled as
    $S_{\nu}=\frac{S_{p}}{1-e^{-1}}\left[1-e^{-(\nu/\nu_{p})^{\alpha_{\mathrm{thin}}-\alpha_{\mathrm{thick}}}}\right]\left(\frac{\nu}{\nu_{p}}\right)^{\alpha_{\mathrm{thick}}}+S_{0}\nu^{\alpha}.$ (4)
Each model was fitted independently, and the best-fit model was determined through Bayesian model selection. RadioSED computes the Bayes factor to compare the likelihoods of different models, selecting the one with the highest posterior probability.
We adopted a standard Gaussian likelihood non-informative priors (log-uniform, linear-uniform), as described in Table 2 of Kerrison et al. (2024), to ensure a statistically robust classification of spectral types that provide a reliable characterisation of the blazar SEDs in our sample. However, the SED modelling does not account for the intrinsic variability between epochs, which can affect the interpretation of blazar spectra. This represents a limitation of our modelling approach that should be considered when analysing the results.
@@ 4 Results
From the SED fitting of the radio emission, we retrieved the spectral behaviour of the sources and found that most of them are well described by the re-triggered peaked spectrum as defined by Equation 4. This smooth behaviour, in contrast to a flat spectrum, suggests that the emission is dominated by a single emitting region rather than by the superposition of multiple compact, self-absorbed regions, as indicated by the classic blazar paradigm (Massardi et al. 2016).
The main parameters retrieved from the SED fitting (Table 2, Figure 4) are the following:
  • $\nu_{\mathrm{trough}}$: This frequency corresponds to the minimum of the SED. After this point, if present, the SED starts to rise.
  • $\alpha_{\mathrm{trough}}$: This is the spectral index at frequencies $\nu<\nu_{\mathrm{trough}}$, which usually corresponds to GLEAM-X photometric points for our sample.
  • $\nu_{\mathrm{peak}}$: This is the frequency where, if present, the peak of the SED takes place.
  • $\alpha_{\mathrm{thick}}$: This is the spectral index at frequencies $\nu_{\mathrm{trough}}<\nu<\nu_{\mathrm{peak}}$, which corresponds to the thick part of the peaked component.
  • $\alpha_{\mathrm{thin}}$: This is the spectral index at frequencies $\nu>\nu_{\mathrm{peak}}$, which corresponds to the thin part of the peaked component.
Using the advanced surveys employed in this study, we conclude that at GLEAM-X wavelengths, the spectral behaviour is not completely flat ($\alpha\sim-0.5$, Table 2). This finding opens the path to an additional extended steep-spectrum emission that was previously missed (Massaro et al. 2013a, b; Nori et al. 2014). GLEAM-X is in fact sensitive to diffuse low-frequency emission likely from older radio lobes that are usually missed in higher-frequency surveys.
Figure 5: Relation between the trough frequency, $\nu_{\text{trough}}$, and the spectral index at lower frequencies, $\alpha_{\text{trough}}$. The colour scale represents the peak frequency, $\nu_{\text{peak}}$, which can be used as a proxy for the cooling time of the source. The colours of the markers’ edges represent the three blazar populations: FSRQs, BL Lacs, and BCUs.
Figure 6: Radio colour-colour diagram of the blazars in our sample based on the spectral indices retrieved from the SED fitting. The colours represent the three blazar populations: FSRQs, BL Lacs, and BCUs.
@@
For most of the sources, the SED extends smoothly towards higher frequencies with a peaked spectrum emission, where it is possible to distinguish a thick part ($\alpha_{\mathrm{thick}}$) and a thin part ($\alpha_{\mathrm{thin}}$). The median $\nu_{\mathrm{peakrest}}$ is $\sim$ 4.6 GHz and no hint of breaks due to ageing is present at high frequency.
In the radio colour-colour diagram, displayed in Figure 6, it emerges that most of the objects ($\sim$ 57.6 %) are in the fourth quadrant; that is, they have a hint of a peaked spectrum, and $\sim$ 10.2% have a flat spectrum after $\nu_{\mathrm{trough}}$. This is a rather small percentage, and it indicates that in the gigahertz domain as well, we should expect a single emitting region rather than the superposition of a number of compact regions of different sizes self-absorbed at different frequencies.
This single component dominating at $\nu>\nu_{\mathrm{trough}}$ could represent the compact self-absorbed segment that was re-activated at a later time. Meanwhile, the low-energy segment of the SED could delineate an extended remnant from earlier activity (Figure 7). Our results suggest that resolution differences have a limited impact on the main trends observed in the radio SEDs, as most sources are unresolved at gigahertz frequencies. Nevertheless, deeper high-resolution, low-frequency surveys (e.g. LOFAR, SKA-Low) will be valuable for disentangling diffuse extended emission from compact jet components.
On the radio colour-colour diagram (Figure 6), BL Lacs display a wide range of behaviours, whereas FSRQs tend to cluster in the flat region. This indicates that FSRQs undergo highly efficient electron pumping, which significantly boosts high-energy radio synchrotron emissions.
Conversely, the two populations seem indistinguishable when considering $\nu_{\mathrm{trough}}$ and $\alpha_{\mathrm{trough}}$ (Figure 6). This may suggest that the extremely low radio frequency emissions are remnants of the source’s history, remaining unchanged by subsequent activities detected at higher frequencies, as the former emissions are confined to the outskirts, while the latter is linked to the nuclear core region.
On the impact of radio variability
Due to the non-simultaneity of the multi-frequency radio data used in this work, some level of spectral distortion may result from intrinsic variability, especially in the most compact and core-dominated blazars. While we did not filter our sample based on observation dates, in many cases the frequency bands were observed within time spans short enough to reduce major inconsistencies (see Section 3). Still, care must be taken in the interpretation of features such as spectral turnovers or curvature, particularly in the higher frequency domain, where variability tends to be more pronounced. In future work, dedicated multi-epoch or strictly simultaneous observations will help constrain the temporal evolution of blazar spectra further and disentangle variability-driven effects from intrinsic spectral shapes.
@@ 5.1 The radio-gamma correlation
Figure 8: Radio luminosity at 150 MHz (left), 855 MHz (centre), and 20 GHz (right) as a function of the $\gamma$-ray luminosity from Fermi-LAT. The black dashed line and the green line indicate the bisector and the median trend, respectively. The colours indicate the blazar class from the 4FGL catalogue.
Table 3: Median and correlation between the ratio of radio and $\gamma$-ray luminosities for different object classes.
All FSRQ BL Lac
@@
The correlation between radio and $\gamma$-ray emission, if present, supports the idea that they are both dominated by the relativistic jet (Giroletti et al. 2016; Mahony et al. 2016). In this work, how this relation behaves in different radio bands can be observed (Figure 8, Table 3). This is particularly useful since the dominant emitting component varies along the radio spectrum, as we found through the SED fitting. The luminosity at 855 MHz is closely tied to the $\gamma$-ray luminosity, indicating that at this frequency, the jet significantly impacts both radio and gamma-ray energy emissions, with a tight relation (correlation coefficient $\rho\sim 0.70$, see Figure 8). This suggests that at 855 MHz, we primarily observe the most powerful segment of the jet nearest the nucleus (Figure 7), which potentially represents the re-activated section. This is supported by the way the radio–gamma-ray relation appears at 20GHz. Consequently, at high radio frequencies, the radio emission is overwhelmingly driven by a strong jet.
In Figure 8 the correlation between 150 MHz and $\gamma$ emission is displayed. Although in this case as well there is a correlation between high- and low-energy emissions (correlation coefficient $\sim 0.70$), there is an evident offset. A potential explanation for this might be that the extended synchrotron emission, which is not efficient enough at generating $\gamma$-rays, dominates this part of the spectrum. This scenario is in full agreement with the results obtained from the SED fitting (Section 4).
The FSRQs dominate the bright end of the luminosity in both radio and $\gamma$, and the radio-gamma correlation holds for FSRQs and BL Lacs. Nevertheless, as shown in Table 3, the radio-gamma correlation assumes higher values for FSRQs at 150 and 855 MHz, which approaches the point at which the radio and $\gamma$ emission are perfectly balanced at lower frequencies. Thus, the FSRQs could already saturate the relation at lower frequencies. This suggests that FSRQs might generally exhibit more effective high-energy pumping across various scales and reflects the presence of a flat spectrum at $\nu>\nu_{trough}$; that is, the emitting mechanism at 855 and 20 GHz is almost the same.
5.2 The radio–X-ray correlation
Figure 9: Radio luminosity at 150 MHz (left), 855 MHz (centre), and 20 GHz (right) as a function of the X-ray luminosity from eROSITA. The black dashed line indicates the bisector. The colours indicate the blazar class from the 4FGL catalogue.
@@ 5.2 The radio–X-ray correlation
Table 4: Median and correlation between the ratio of radio and X-ray luminosities for different object classes.
All FSRQ BL Lac
@@
Correlation 0.82 0.48 0.99
The main driver of the emission in this range is the hot corona, which closely traces the accretion disc, especially in FSRQs. The radio–X-ray correlation is shown in Figure 9. We show the results in Table 4, and we found almost no correlation for the 855 MHz but a tight one at 150 MHz. In this particular situation, when even giving just a glance at Figure 9, it is clear that the global behaviour might be misleading and not truly representative of the physical processes taking place. The real picture emerges only when distinguishing between different classes of blazars. In FSRQs, radio and soft X-ray emissions are tightly correlated, while in BL Lacs the relationship is much more scattered.
The results we obtained (Table 4) confirm that FSRQs are in an accretion-driven regime, which fuels both the radio jet and the disc emission. On the other hand, in BL Lacs, the jets are mainly rotation driven. However, alternative factors could contribute to the observed trends. For instance, BL Lacs exhibit a higher variability in the X-ray band, which may introduce additional scatter in the correlation (Ghisellini et al. 2017).
Moreover, the impact of the circumnuclear medium could play a role in shaping the X-ray emission properties, particularly in FSRQs, where disc-related processes contribute significantly to the total luminosity. A deeper understanding of these effects will require time domain studies incorporating simultaneous X-ray and radio observations.
When comparing the X-ray luminosity with the 20 GHz emission, it is essential to account for selection bias. The sources in our sample that have both a 20 GHz and an X-ray counterpart represent a highly efficient subset of our data. Consequently, the tighter correlation between X-ray and radio emissions is not surprising given that the fraction of FSRQs is higher at these frequencies compared to lower frequencies. Deeper millimetre observations are required to gain a comprehensive scientific understanding of the behaviour at higher frequencies.